# Hybrid Quantum Solvers in Production: how to succeed in the NISQ era?

### 25 Jan 2024 v3

@@ Abstract
Hybrid quantum computing is considered the present and the future within the field of quantum computing. Far from being a passing fad, this trend cannot be considered just a stopgap to address the limitations of NISQ-era devices. The foundations linking both computing paradigms will remain robust over time. Despite buoyant research activity, the challenges in hybrid computing are still countless, ranging from the proper characterization of current solvers to the establishment of appropriate methodologies for the design and fair evaluation of hybrid algorithms. The contribution of this work is twofold: first, we describe and categorize some of the most frequently used hybrid solvers, resorting to two different taxonomies recently published in the literature. Secondly, we put a special focus on two solvers that are currently deployed in real production and that have demonstrated to be near the real industry. These solvers are the LeapHybridBQMSampler contained in D-Wave’s Hybrid Solver Service and Quantagonia’s Hybrid Solver. We analyze the performance of both hybrid methods using as benchmarks four well-known combinatorial optimization problems: the Traveling Salesman Problem, Vehicle Routing Problem, Bin Packing Problem, and Maximum Cut Problem. Thanks to the contributions presented in this paper, the reader gains insight into the performance of those hybridization strategies nowadays in production and close to the industrial markets.
@@ I Introduction
Furthermore, it would be a mistake to conceive of quantum-classical hybrid computing simply as a stopgap to address the limitations of NISQ-era computers, just as quantum computing should not be referred to as a natural substitute for classical computing. As stated in [3], ”hybrid algorithms are likely here to stay well past the NISQ era and even into full fault-tolerance, with the quantum processors augmenting the already powerful classical processors which exist by performing specialized task”. In other words, classical and quantum computing will play collaborative and cooperative (but never competitive) roles in the mid-term horizon.
@@
      According to the layout established by the workflow of the hybrid solver, following the guidelines of [4]. This means focusing on the logical arrangement of the quantum and classical modules within the method. Note that this work focuses on giving an overview of the hybrid solver strategies with emphasis on the designable-by-designer steps. This means that the taxonomy will not include any consideration with respect to vertical hybrid workflows, i.e. those with ”controlling activities required to control and operate” either a classical or a quantum subroutine, as these operations are not key functional stages of the algorithmic scheme of the solver.
@@
  • Given the previous theoretical framework, gaining insight into the performance of those hybridization strategies nowadays in production and close to industrial markets. To do that, an experimentation over two hybrid solvers is conducted: the LeapHybridBQMSampler (LeapBQM) contained in D-Wave’s Hybrid Solver Service (HSS, [6]), and Quantagonia’s Hybrid Solver [^1] https://www.quantagonia.com/hybridsolver. For these tests, the QOPTLib benchmark [7] has been employed, which is composed of 40 instances equally distributed over four well-known combinatorial optimization problems: the Traveling Salesman Problem (TSP, [8]), the Vehicle Routing Problem (VRP, [9]), the one-dimensional Bin Packing Problem (BPP, [10]), and the Maximum Cut Problem (MCP, [11]).
@@
The rest of this paper is structured as follows: In the following Section II, we perform the classification of five of the most renowned and used hybrid solvers to finally sum up with the current trends for the hybrid solver designs which answers the main question posed in the title of the section: ”Where might the quantum advantage be?”. After that, Section III describes the experimentation conducted with LeapBQM and QHS, including the experimental setup and results. Finally, Section IV closes this document by drawing some conclusions and outlining future research lines.
@@ II Classifying hybrid solvers: Where might the quantum advantage be?
However, there are a number of perfectly recognizable schemes that have been frequently employed by the community to deal with optimization problems. From the perspective of the work in [4], all of these configurations are horizontal hybrid schemes since they are pipelines containing ”all operational activities required to use a quantum computer and a classical computer to perform an algorithm”.
@@
  • Variational Quantum Algorithms (VQAs): VQAs are the most used hybrid schemes for solving optimization problems through the gate-based quantum paradigm. As briefly explained in [15], ”the trademark of VQAs is that they use a quantum computer to estimate the cost function of a problem (or its gradient) while leveraging the power of classical optimizers to train the parameters of the quantum circuit”. The most representative examples of the VQA framework are the Variational Quantum Eigensolver (VQE, [16]) and the Quantum Approximate Optimization Algorithm (QAOA, [17]). We refer readers interested in VQE and QAOA applications to [18] and [19].
    Following the taxonomy presented in [5] and represented in Figure 1, VQAs can be considered as hybrid schemes with a classical method optimizing the circuit parameters, thus playing a cooperative role where the solver ”contains a core intelligence engine composed of both classical and quantum artifacts”. Furthermore, in VQAs, there is no clear division of exploration and exploitation activities.
@@
  • QBSolv: QBSolv is a hybrid method developed by D-Wave, which can be described as ”a decomposing solver for solving a QUBO problem by splitting it into pieces [^2] https://github.com/dwavesystems/qbsolv. The decomposing procedure of this method starts with an initial solution provided by a classical Tabu Search algorithm [20]. A quantum method later improves that solution by sending a user-specified fraction of it to the QPU. This subproblem includes the most energetic variables, of size fraction, of the initial problem. In practice, this value is usually between 0.05 and 0.15.
    Despite this technique was deprecated as of the end of 2021 and discontinued after March 2022, QBSolv has been widely used by researchers, even in recent studies [21, 22, 23]. Embracing the Figure 1 taxonomy, the quantum and classical procedures pull together with an imbricated, hence cooperative, scheme where ”both participate in searching for the solution of the problem[5]. Figure 3 displays the general workflow of QBSolv.
@@
One step forward from these standard hybrid schemes is taken when introducing parallel computing in advanced hybrid solver designs. Accordingly and in conformity with the criteria set out in [4], the following solvers fall within the parallel hybrid category since ”the workflows of parallel hybrid have multiple independent branches to solve a specific problem”. Arguably, the most representative example is the D-Wave-Hybrid-Framework [24].
As explained by its creators, the D-Wave-Hybrid-Framework is ”a general, minimal Python framework for building hybrid asynchronous decomposition samplers for QUBO problems [^3] https://docs.ocean.dwavesys.com/en/stable/docs_hybrid/. This framework is appropriate for ”developing hybrid approaches to combining quantum and classical compute resources”. Figure 4 represents the general scheme of the D-Wave-Hybrid-Framework as it was conceived in [24]. The best known concretization is the solver coined as Kerberos. This solver is a reference hybrid workflow which is composed of three different techniques running in parallel for a number of iterations: a quantum one that accesses the QPU and two classical ones, a Tabu and a Simulated Annealing [25]. It should be noted that these branches share solutions along the execution. That is, the best solution found by the branches in an iteration is fed into all the branches in the next iteration. Recent practical applications of Kerberos can be found in papers such as [26, 27].
@@
      Next, HSS generates one or more hybrid threads devoted to solving the problem at hand. Each thread counts with a Classical Heuristic Module (CH), which is in charge of exploring the complete solution space, and a Quantum Module (QM), which, in the words of D-Wave, ”formulates quantum queries that are sent to a back-end Advantage QPU. Replies from the QPU are used to guide the heuristic module toward more promising areas of the search space or to find improvements to existing solutions”. In other words, QM is mainly in charge of exploitation duties. Finally, the solver forwards to the user the best solution among those found in the pool of threads.
@@
    Given this definition, HSS belongs to the the cooperative hybrid category, and more concretely to the imbricated schemes.
  • Quantagonia’s Hybrid Solver: established in November 2021, Quantagonia’s main objective is to bridging the gap between quantum computing, today mainly in lab, and real industry needs. Among the portfolio of products offered by this German company, QHS is a hardware-agnostic method for addressing Linear Programming, Mixed-Integer Programming and QUBO problems. Despite that QHS has not been employed in any research paper yet, it is currently being used by leading companies such as Strangeworks, MathPlan, T-Systems, and Adesso, which reinforces its practical application.
@@
    Furthermore, and being one of its strengths, QHS includes an optimality proof mechanism, which means that, along with the results, the solver also provides the optimality gap. The optimality gap tells how much more potential there could be in the optimization process. This gap would be 0% whether the solution found is proven to be the optimal one. Another advantage of QHS in contrast with HSS lies in its flexibility and ample configurability. Despite being a solver conceived to be in production, QHS gives a greater control to the user, such as the (de)activation of the Branch-and-Bound step, the addition of further primal heuristics, or the setting of a time limit, among many others. On the contrary, HSS methods are blackbox algorithms just allowing a maximum runtime to be set.
@@
Figure 6: Quantagonia’s Hybrid Solver workflow. CH = classical heuristic. QM = Quantum Module. B&B = Branch and Bound algorithm.
@@
VQA Cooperative - Parameter Optimization
Responsability shared
QBSolv Cooperative - Imbrication
Exploration: classical computing, Exploitation: QPU,
D-Wave-Hybrid- Framework Cooperative - Imbrication
Responsability shared
HSS
Cooperative - Imbrication (each generated thread)
Exploration: classical computing, Exploitation: QPU,
QHS Cooperative - Imbrication
Exploration: QPU, Exploitation: classical computing,
@@ III Experimentation
We refer interested readers to [7] for gathering further information on the QOPTLib.
@@
Thus, Table II shows the results obtained. For each instance, we represent the result average obtained for each solver, as well as the standard deviation and the median value. Note that all these values correspond to the original objective functions above described. Aiming to enhance the replicability of this work, all the 40 instances used and the results obtained are openly available in [33].

### 19 Mar 2024 v4

@@ Abstract
Hybrid quantum computing is considered the present and the future within the field of quantum computing. Far from being a passing fad, this trend cannot be considered just a stopgap to address the limitations of NISQ-era devices. The foundations linking both computing paradigms will remain robust over time. The contribution of this work is twofold: first, we describe and categorize some of the most frequently used hybrid solvers, resorting to two different taxonomies recently published in the literature. Secondly, we put a special focus on two solvers that are currently deployed in real production and that have demonstrated to be near the real industry. These solvers are the LeapHybridBQMSampler contained in D-Wave’s Hybrid Solver Service and Quantagonia’s Hybrid Solver. We analyze the performance of both methods using as benchmarks four combinatorial optimization problems.
@@ I Introduction
Furthermore, it would be a mistake to conceive of quantum-classical hybrid computing simply as a stopgap to address the limitations of NISQ-era computers, just as quantum computing should not be referred to as a natural substitute for classical computing. As stated in [3], “hybrid algorithms are likely here to stay well past the NISQ era and even into full fault-tolerance, with the quantum processors augmenting the already powerful classical processors which exist by performing specialized task”. In other words, classical and quantum computing will play collaborative and cooperative (but never competitive) roles in the mid-term horizon.
@@
      According to the layout established by the workflow of the hybrid solver, following the guidelines of [4]. This means focusing on the logical arrangement of the quantum and classical modules within the method. Note that this work focuses on giving an overview of the hybrid solver strategies with emphasis on the designable-by-designer steps. This means that the taxonomy will not include any consideration with respect to vertical hybrid workflows, i.e. those with “controlling activities required to control and operate” either a classical or a quantum subroutine, as these operations are not key functional stages of the algorithmic scheme of the solver.
@@
  • Given the previous theoretical framework, gaining insight into the performance of those hybridization strategies nowadays in production and close to industrial markets. To do that, an experimentation on two hybrid solvers is conducted: the LeapHybridBQMSampler (LeapBQM) contained in D-Wave’s Hybrid Solver Service (HSS, [6]), and Quantagonia’s Hybrid Solver [^1] https://www.quantagonia.com/hybridsolver. For these tests, the QOPTLib benchmark [7] has been employed, which is composed of 40 instances equally distributed over four combinatorial optimization problems: the Traveling Salesman Problem (TSP, [8]), the Vehicle Routing Problem (VRP, [9]), the one-dimensional Bin Packing Problem (BPP, [10]), and the Maximum Cut Problem (MCP, [11]).
@@
The rest of this paper is structured as follows: In the following Section II, we perform the classification of five of the most renowned and used hybrid solvers to summarize current trends for hybrid solver designs, which answers the main question posed in the title of the section: “Where might the quantum advantage be?”. After that, Section III describes the experimentation conducted with LeapBQM and QHS, including the experimental setup and results. Finally, Section IV closes this document by drawing some conclusions and outlining future research lines.
@@ II Classifying hybrid solvers: Where might the quantum advantage be?
However, there are a number of perfectly recognizable schemes that have been frequently employed by the community to deal with optimization problems. From the perspective of the work in [4], all of these configurations are horizontal hybrid schemes since they are pipelines containing “all operational activities required to use a quantum computer and a classical computer to perform an algorithm”.
@@
  • Variational Quantum Algorithms (VQAs): VQAs are the most used hybrid schemes for solving optimization problems through the gate-based quantum paradigm. As briefly explained in [15], “the trademark of VQAs is that they use a quantum computer to estimate the cost function of a problem (or its gradient) while leveraging the power of classical optimizers to train the parameters of the quantum circuit”. The most representative examples of the VQA framework are the Variational Quantum Eigensolver (VQE, [16]) and the Quantum Approximate Optimization Algorithm (QAOA, [17]). We refer readers interested in VQE and QAOA applications to [18] and [19].
    Following the taxonomy presented in [5] and represented in Figure 1, VQAs can be considered as hybrid schemes with a classical method optimizing the circuit parameters, thus playing a cooperative role where the solver “contains a core intelligence engine composed of both classical and quantum artifacts”. Furthermore, in VQAs, there is no clear division of exploration and exploitation activities.
@@
  • QBSolv: QBSolv is a hybrid method developed by D-Wave, which can be described as “a decomposing solver for solving a QUBO problem by splitting it into pieces [^2] https://github.com/dwavesystems/qbsolv. The decomposing procedure of this method starts with an initial solution provided by a classical Tabu Search algorithm [20]. A quantum method later improves that solution by sending a user-specified fraction of it to the QPU. This subproblem includes the most energetic variables, of size fraction, of the initial problem. In practice, this value is usually between 0.05 and 0.15.
    Despite the fact this technique was deprecated as of the end of 2021 and discontinued after March 2022, QBSolv has been widely used by researchers, even in recent studies [21, 22, 23]. Embracing the Figure 1 taxonomy, the quantum and classical procedures pull together with an imbricated, hence cooperative, scheme where “both participate in searching for the solution of the problem[5]. Figure 3 displays the general workflow of QBSolv.
@@
One step forward from these standard hybrid schemes is taken when introducing parallel computing in advanced hybrid solver designs. Accordingly and in conformity with the criteria set out in [4], the following solvers fall within the parallel hybrid category since “the workflows of parallel hybrid have multiple independent branches to solve a specific problem”. Arguably, the most representative example is the D-Wave-Hybrid-Framework [24].
As explained by its creators, the D-Wave-Hybrid-Framework is “a general, minimal Python framework for building hybrid asynchronous decomposition samplers for QUBO problems [^3] https://docs.ocean.dwavesys.com/en/stable/docs_hybrid/. This framework is appropriate for “developing hybrid approaches to combining quantum and classical compute resources”. Figure 4 represents the general scheme of the D-Wave-Hybrid-Framework as it was conceived in [24]. The best known concretization is the solver coined as Kerberos. This solver is a reference hybrid workflow which is composed of three different techniques running in parallel for a number of iterations: a quantum one that accesses the QPU and two classical ones, a Tabu and a Simulated Annealing [25]. It should be noted that these branches share solutions along the execution. That is, the best solution found by the branches in an iteration is fed into all the branches in the next iteration. Recent practical applications of Kerberos can be found in papers such as [26, 27].
@@
      Next, HSS generates one or more hybrid threads devoted to solving the problem at hand. Each thread counts with a Classical Heuristic Module (CH), which is in charge of exploring the complete solution space, and a Quantum Module (QM), which, in the words of D-Wave, “formulates quantum queries that are sent to a back-end Advantage QPU. Replies from the QPU are used to guide the heuristic module toward more promising areas of the search space or to find improvements to existing solutions”. In other words, QM is mainly in charge of exploitation duties. Finally, the solver forwards to the user the best solution among those found in the pool of threads.
@@
    Given this definition, HSS belongs to the cooperative hybrid category, and more concretely to the imbricated schemes.
  • Quantagonia’s Hybrid Solver: established in November 2021, Quantagonia’s main objective is to bridging the gap between quantum computing, today mainly in lab, and real industry needs. Among the portfolio of products offered by this German company, QHS is a hardware-agnostic method for addressing Linear Programming, Mixed-Integer Programming and QUBO problems. Despite the fact that QHS has not been employed in any research paper yet, it is currently being used by leading companies such as Strangeworks, MathPlan, T-Systems, and Adesso, which reinforces its practical application.
@@
    Furthermore, and being one of its strengths, QHS includes an optimality proof mechanism, which means that, along with the results, the solver also provides the optimality gap. The optimality gap tells how much more potential there could be in the optimization process. This gap would be 0% when the solution found is proven to be the optimal one. Another advantage of QHS in contrast with HSS lies in its flexibility and ample configurability. Despite being a solver conceived to be in production, QHS gives a greater control to the user, such as the (de)activation of the Branch-and-Bound step, the addition of further primal heuristics, or the setting of a time limit, among many others. On the contrary, HSS methods are blackbox algorithms just allowing a maximum runtime to be set.
@@
Figure 6: Quantagonia’s Hybrid Solver workflow. CH = classical heuristic.
QM = Quantum Module. B&B = Branch and Bound algorithm.
@@
VQA Cooperative - Parameter Optimization Responsibility shared
QBSolv Cooperative - Imbrication Exploration: classical computing, Exploitation: QPU,
D-Wave-Hybrid- Framework Cooperative - Imbrication Responsibility shared
HSS Cooperative - Imbrication (each generated thread) Exploration: classical computing, Exploitation: QPU,
QHS Cooperative - Imbrication Exploration: QPU, Exploitation: classical computing,
@@ III Experimentation
Thus, Table II shows the results obtained after 10 independent runs per instance. For each instance, we represent the result average obtained for each solver, as well as the standard deviation and the median value. Note that all these values correspond to the original objective functions above described. Aiming to enhance the replicability of this work, all the 40 instances used and the results obtained are openly available in [33].